40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 14 20:42:04 2021, gautam, Update, LSC, Locking 
    Reply  Thu Jul 15 14:25:48 2021, Paco, Update, LSC, CM board high_BW_TFs.pdf
Message ID: 16247     Entry time: Wed Jul 14 20:42:04 2021     Reply to this: 16248
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking 

[paco, gautam]

we decided to give the PRFPMI lock a go early-ish. Summary of findings today eve:

  1. Arms under ALS control display normal noise and loop UGFs.
  2. PRMI took longer than usual to lock (when arms are held off resonance) - could be elevated sesimic, but warrants measuring PRMI loop TFs to rule out any funkiness. MICH loop also displayed some saturation on acquisition, but after the boosts and other filters were turned on, the lock seemed robust and the in-loop noise was at the usual levels.
  3. We are gonna do the high bandwidth single arm locking experiments during daytime to rule out any issues with the CM board.

The ALS--> IR CARM handoff is the problematic step. In the past, getting over this hump has just required some systematic loop TF measurements / gain slider readjustments. We will do this in the next few days. I don't think the ALS noise is any higher than it used to be, and I could do the direct handoff as recently as March, so probably something minor has changed.

ELOG V3.1.3-